OpenStudy (hitaro9):

Water leaking onto a floor creates a circular pool with an area that increases at the rate of 3 square centimeters per minute. How fast is the radius of the pool increasing when the radius r is 3 centimeters?

OpenStudy (hitaro9):

So, I tried doing\[dA/dt = 3\] and \[dA/dt = 2\pi r(dr/dt)\] to get \[(dr/dt) =1/2\pi \] but that didn't work. help please?

zepdrix (zepdrix):

\[\Large A'=3\]And they want us to find \(\Large r'\) when \(\Large r=3\). Your work looks good.\[\Large A=\pi r^2\]\[\Large A'=2\pi r r'\]\[\Large r=\frac{A'}{r' 2 \pi}=\frac{1}{2\pi}\]

zepdrix (zepdrix):

Are you entering this online? Does it require you to enter it like this maybe? D: 1/(2pi)

zepdrix (zepdrix):

Woops I solved for r instead of r' lol :) silly mistake there. Should work out the same though.

OpenStudy (hitaro9):

Ah yeah. I see. I was supposed to enter it in as litterally 1/2pi rather than 1.57. Thanks.